Use the following two equations: QS = 972 + 103.5P QD = 1722 – 141.5P A- Calculate the equilibrium price and the equilibrium quantity. Show all your work. B- Using the above two equations to find the values of Qd, Qs, the market situation (Shortage/Surplus/Equilibrium), and the Value of shortage or surplus if any, at the following prices: 3.044, 3.65, 3.88, and 3.95. C- If the consumer income increases by 40%, what will happen to the equilibrium price and quantity.


1
Expert's answer
2022-05-25T15:52:04-0400

"Q_s=972+103.5P,"

"Q_d=1722-141.5P,"

"Q_s=Q_d,"

"972+103.5P=1722-141.5P,"

"750=245P,"

"P=3.06,"

"Q_s=Q_d=1288.8."